### 18 Dec 2023 v1

@@ Abstract
Seizure events may manifest as transient disruptions in movement and behavior, and the analysis of these clinical signs, referred to as semiology, is subject to observer variations when specialists evaluate video-recorded events in the clinical setting. To enhance the accuracy and consistency of evaluations, computer-aided video analysis of seizures has emerged as a natural avenue. In the field of medical applications, deep learning and computer vision approaches have driven substantial advancements. Historically, these approaches have been used for disease detection, classification, and prediction using diagnostic data; however, there has been limited exploration of their application in evaluating video-based motion detection in the clinical epileptology setting. While vision-based technologies do not aim to replace clinical expertise, they can significantly contribute to medical decision-making and patient care by providing quantitative evidence and decision support. Behavior monitoring tools offer several advantages such as providing objective information, detecting challenging-to-observe events, reducing documentation efforts, and extending assessment capabilities to areas with limited expertise. The main applications of these could be (1) improved seizure detection methods; (2) refined semiology analysis for predicting seizure type and cerebral localization. In this paper, we detail the foundation technologies used in vision-based systems in the analysis of seizure videos, highlighting their success in semiology detection and analysis, focusing on work published in the last 7 years. We systematically present these methods and indicate how the adoption of deep learning for the analysis of video recordings of seizures could be approached. Additionally, we illustrate how existing technologies can be interconnected through an integrated system for video-based semiology analysis. Each module can be customized and improved by adapting more accurate and robust deep learning approaches as these evolve. Finally, we discuss challenges and research directions for future studies.
@@ 1 Introduction
Seizure semiology involves the analysis of stereotypical motor manifestations such as convulsions, tonic, clonic, and hyperkinetic changes which are a source of major disability for patients. Alongside electrophysiological and neuroimaging recordings, the insights derived from semiology are essential, particularly for patients resistant to drug treatment [32, 195, 12]. The clinical expression of seizures represents the primary symptomatic challenge in epilepsy. Seizure semiology, though a crucial data source for clinicians, poses substantial methodological challenges for research, such as the selection of semiologic categories. A more profound understanding of semiologic categorization holds relevance for refining seizure classification systems [132, 139] The interpretation of video monitoring recordings, however, relies on clinical experience and can vary between clinicians and across patients, especially when seizures involve complex signs including hyperkinetic motor behavior [176]. To address this, automated quantification of semiology shows promise for extracting more objective information from these recordings [4, 106, 89]. Quantified analysis of seizures through video is desirable for its non-invasiveness and potentially easy deployment, utilizing existing video hardware in various settings, once clinically relevant algorithms have been developed. A main motivation in developing this field has been the goal of accurate automated seizure detection, because of the potential impact on morbidity, mortality and quality of life in people with epilepsy [113]. Automated methods for seizure detection and classification may also help improve patient monitoring, reducing the time and effort needed for screening long-term video-EEG data in specialized epilepsy monitoring units (EMUs). Better methods for automated video-based seizure detection can also be applied to home recording systems, once validated in a hospital setting [158]. Lastly, as in other clinical fields, artificial-intelligence-enhanced diagnostic methods may be trained to “see” patterns that clinicians do not reliably recognize, with the potential for better prediction of cerebral localization from complex semiologic features, such as those seen in hyperkinetic seizures. This last application in particular is under-developed due to challenges such as the scarcity of datasets and the complexity of the natural clinical setting, which need to be addressed in this field. The goal of the present review is to provide an overview of advances in the field, focusing on work published from 2016 onwards (date of our previous review), and highlighting ongoing challenges and future directions.
@@ 1.1 Vision-based action recognition in health
Machine learning has significantly influenced various technological domains such as signal and human motion analysis, especially with the advent of modern deep learning approaches. Deep learning, which constitutes a subset of machine learning, involves advanced neural networks that enable a system to automatically learn and understand hierarchical features, providing a more sophisticated approach to interpreting complex data [117]. This impact is seen in the field of video-based human action analysis, which has become a focal point in the field of computer vision and pattern recognition. Recent technical advances have allowed increasingly detailed analyses of action, gestures, and facial expressions [217]. Human action recognition, a key aspect of semiology analysis, aims to identify different actions from sequences of observations across diverse environmental conditions. The applications of such recognition methods are wide-ranging, contributing to advancements relating to human health and performance. Applications include assessment of human development, optimization of human performance, neuromuscular rehabilitation, prevention of musculoskeletal injury, gait analysis, and motor assessment [188, 19]. Existing literature extensively explores the causal relationship between diseases and abnormal body actions such as atypical head poses and movements, emotions, eye movements, as well as upper and lower limb movements [52, 199].
1.2 Vision-based seizure analysis
@@
There is an increasing interest in leveraging video analysis techniques for the automated analysis of movements associated with motor stereotypes such as Parkinson’s disease, tremors, peripheral neuropathy, Huntington’s disease, ataxia, dystonia, and Epilepsy [36, 142, 146, 95]. Vision-based techniques, in particular, have proven valuable in patient monitoring [174]. While general vision-based motion analyses excel in uncluttered settings, they can be unreliable in noisy environments like epilepsy monitoring and intensive care units. These settings pose unique challenges, including varying lighting conditions throughout recording sessions, environmental occlusions (e.g. bed blankets, head wrapping, hospital gown), and interference from non-subject entities (e.g. clinician, nurse, visitor) [196]. While current state-of-the-art deep learning models exhibit promise in addressing these challenges in epilepsy, their effectiveness is still in the early stages. We can anticipate that more advanced video-based automatic semiologic categorization will become possible as the field progresses. Existing methods struggle to recognize subject-specific semiologic categories of unseen behaviors, yielding modest performance. Achieving fine-grained semiology recognition, crucial for distinguishing the stepwise progression of clinical features, remains a challenge. In addition, ensuring effective explainability to understand model behavior beyond traditional performance indicators is essential, with a particular focus on clinician usability. As these translations of research progress, there is a need for research to facilitate the proper clinical adoption of methods, and to build trust between clinical experts and AI frameworks.
1.3 Contributions of the present review
@@
As the field is rapidly evolving and researchers are interested in understanding the strengths and pitfalls of deep learning-based solutions for the analysis of seizure semiology, a comprehensive survey is required that provides insights from existing works and motivates researchers to address limitations and challenges within this field. Conventional motion analysis of seizures, employing classical markerless computer vision methods were reviewed by Pediaditis et al. [157] and Abbasi et al. [1]. A more detailed examination, considering data-driven solutions for motion capture and action detection, was presented by Ahmedt-Aristizabal et al. [7]. A recent work by Karacsony et al. [103] provided insights into advancements in video-based clinical in-bed monitoring, particularly within EMUs. The work discusses opportunities to address challenges associated with 3D motion capture, including domain transfer, occlusions, varying viewpoints, and low resolution. In contrast to other reviews that focus on theoretical aspects and applications of action recognition in semiology, our manuscript makes novel contributions, summarized as follows:
@@
  • We extend our 2016 review of vision-based approaches for semiology analysis [7] by incorporating all relevant works concerning the adoption of deep learning techniques to enhance the automated detection and quantification of seizures, and/or of individual seizure features captured on video recordings. All deep learning methods adopted by these works are described in detail and categorized according to the motion analysis framework used. Case studies reported in this paper are obtained from various journals, conference proceedings, and open-access repositories published in English between 2016 and 2023.
@@
  • Our paper identifies several challenges specific to action recognition models in the context of seizure semiology based on video analysis, which have not been comprehensively addressed in previous reviews. These challenges cover areas such as motion capture of stereotyped movements and domain adaptation, privacy preservation, interpretation of action models, open-set action analysis, and time-evolving or fine-grain action recognition. Existing research efforts in AI and computer vision addressing these challenges are highlighted.
@@
This manuscript comprises three main sections. In Section 2, we provide a technical overview of current approaches for vision-based human action recognition (HAR) in Epilepsy. Specifically, Section 2.3, introduces skeleton-based HAR approaches, while Section 2.4 focuses on spatio-temporal-based (frame-based) HAR models. In Section 3, we introduce the current applications of computer vision and deep learning techniques for semiology analysis. Details regarding the existing approaches along with the computer vision techniques are categorized according to the motion framework employed. Section 4 offers guidance on incorporating existing approaches along with the computer vision techniques they employ for clinical use through a multi-stream framework for semiology analysis. Finally, in Section 5, we clarify limitations and offer new perspectives on the current state of semiology-based seizure analysis.
@@ 2.1 Overview of modern deep learning and human action recognition
Deep learning, a subset of machine learning, enables the creation of computational architectures comprising multiple processing layers. These architectures learn data representations with varying levels of abstraction through non-linear transformations [27]. Emulating the operation of the human brain, deep learning methods utilize multiple hierarchical layers alongside sophisticated training algorithms. The primary distinction from traditional machine learning lies in feature engineering. In deep learning algorithms, feature representations are automatically derived from the training data rather than relying on human assumptions as seen in traditional hand-crafted approaches [117]. This proves particularly beneficial in scenarios where understanding the relationship between input and output is limited [15]. Deep learning techniques have become predominant in computer vision, outperforming traditional approaches across various tasks, including human action recognition. Noteworthy deep neural network architectures include convolutional neural networks (CNNs), recurrent neural networks (RNNs), graph neural networks (GNNs) and the Transformer model.

### 04 Mar 2024 v2

@@ Abstract
Seizure events can manifest as transient disruptions in the control of movements which may be organized in distinct behavioral sequences, accompanied or not by other observable features such as altered facial expressions. The analysis of these clinical signs, referred to as semiology, is subject to observer variations when specialists evaluate video-recorded events in the clinical setting. To enhance the accuracy and consistency of evaluations, computer-aided video analysis of seizures has emerged as a natural avenue. In the field of medical applications, deep learning and computer vision approaches have driven substantial advancements. Historically, these approaches have been used for disease detection, classification, and prediction using diagnostic data; however, there has been limited exploration of their application in evaluating video-based motion detection in the clinical epileptology setting. While vision-based technologies do not aim to replace clinical expertise, they can significantly contribute to medical decision-making and patient care by providing quantitative evidence and decision support. Behavior monitoring tools offer several advantages such as providing objective information, detecting challenging-to-observe events, reducing documentation efforts, and extending assessment capabilities to areas with limited expertise. The main applications of these could be (1) improved seizure detection methods; (2) refined semiology analysis for predicting seizure type and cerebral localization. In this paper, we detail the foundation technologies used in vision-based systems in the analysis of seizure videos, highlighting their success in semiology detection and analysis, focusing on work published in the last 7 years. We systematically present these methods and indicate how the adoption of deep learning for the analysis of video recordings of seizures could be approached. Additionally, we illustrate how existing technologies can be interconnected through an integrated system for video-based semiology analysis. Each module can be customized and improved by adapting more accurate and robust deep learning approaches as these evolve. Finally, we discuss challenges and research directions for future studies.
@@ 1 Introduction
Seizure semiology involves the analysis of seizure clinical manifestations, which may include stereotypical behavioral and motor changes, among them tonic, clonic or complex motor patterns. The clinical expression of seizure represents the primary symptomatic challenge in epilepsy. Various semiologic manifestations of seizures are potentially measurable using video analysis: motor alterations (e.g. as seen in automatic behaviors, hyperkinetic seizures, or tonic-clonic seizures); an abnormal lack of movement (typical of some seizures with prominent loss of responsiveness to the environment); or an abrupt change in posture that constitutes a risk factor for falls (e.g. in atonic seizures). Alongside electrophysiological recordings and neuroimaging, the insights derived from semiology are essential, particularly for patients resistant to drug treatment [35, 218, 12]. Seizure semiology, though a crucial data source for clinicians, poses substantial methodological challenges for research, such as the selection of semiologic categories. A more profound understanding of semiologic categorization holds relevance for refining seizure classification systems [152, 159]. The interpretation of video monitoring recordings, however, relies on clinical experience and can vary between clinicians and across patients, especially when seizures involve complex signs including hyperkinetic motor behavior [197].
Therefore automated quantification of semiology shows promise for extracting more objective information from these recordings [4, 122, 105]. Quantified analysis of seizures through video is desirable for its non-invasiveness and potentially easy deployment, utilizing existing video hardware in various settings, once clinically relevant algorithms have been developed. A main motivation in developing this field has been the goal of accurate automated seizure detection, because of the potential impact on morbidity, mortality and quality of life in people with epilepsy [132].
Expanding on the interdisciplinary nature of this research, it is noteworthy that although the characterization of behavior, in general, needs to be placed in the context of biology and specifically of ethology, the systematic study of behavior [221] has evolved into quantitative methods such as neuroethology [77] and computational neuroethology [58]. These concepts, associated with the study of the complexity of behavioral sequences such as those present in human epilepsies [89, 56, 30, 216], offer a unique perspective. The application of these approaches to epilepsy research allows for a more integrated understanding of the complex interactions between neural circuits and behavior during seizure activity. Moreover, they provide a solid foundation for developing computational models that bridge the gap between clinical observation and automated quantification.
@@
Automated methods for seizure detection and classification may also help improve patient monitoring, reducing the time and effort needed for screening long-term video-EEG data in specialized Epilepsy monitoring units (EMUs). Better methods for automated video-based seizure detection can also be applied to home recording systems, once validated in a hospital setting [178]. Lastly, as in other clinical fields, artificial-intelligence-enhanced diagnostic methods may be trained to “see” patterns that clinicians do not reliably recognize, with the potential for better prediction of cerebral localization from complex semiologic features, such as those seen in hyperkinetic seizures. This last application in particular is under-developed due to challenges such as the scarcity of datasets and the complexity of the natural clinical setting, which need to be addressed in this field.
1.1 What is modern deep learning and computer vision?
The growing use of artificial intelligence (AI) and machine learning has significantly influenced various technological domains, such as signal and human motion analysis. Machine learning, a subset of AI, encompasses diverse methods that enable systems to automatically learn patterns and make predictions from data. The application of these learning methods, particularly in the domain of video analysis, has gained prominence with the advent of modern deep learning approaches. Deep learning, a subset of machine learning, employs many-layered neural networks to automatically learn hierarchical features from data [137]. Notably, the rise of deep learning has seen the field of video-based human action analysis become a focal point within the field of computer vision and pattern recognition, with recent technical advances allowing increasingly detailed analyses of actions, gestures, and facial expressions [243].
A foundational learning paradigm in machine learning is supervised learning, where an algorithm is trained on a labeled dataset, with input data and the corresponding desired outputs provided. From this, a model is trained to learn the mapping between inputs and outputs. As such techniques are data-driven and dependent on the quantity and quality of the provided data, there are several critical challenges for medical applications including scarce annotation (too little labelled data), complexity (a single type of instance to detect may appear in many different forms), weak annotations (labels may be noise), and label sparsity (some classes have very few instances). To address these challenges, several other learning paradigms have been proposed in the literature. Weakly supervised learning emerges as a strategy, leveraging partially annotated examples to navigate the scarcity of labeled data. Self-supervised learning adopts a different approach, where the data itself provides supervisory signals, aiding in learning representations through proxy tasks. Additionally, reinforcement learning introduces a concept where an agent learns optimal behavior by interacting within a dynamic environment. Further background concepts in the domain of computer vision and deep learning are discussed in Section 3.
@@ 1.2 Vision-based action recognition in health
Human action recognition, a key aspect of semiology analysis, aims to identify different actions from sequences of observations across diverse environmental conditions. The applications of such recognition methods driven by deep learning and computer vision are wide-ranging, contributing to advancements relating to human health and performance. Applications include assessment of human development, optimization of human performance, neuromuscular rehabilitation, prevention of musculoskeletal injury, gait analysis, and motor assessment [210, 19]. Existing literature extensively explores the causal relationship between diseases and abnormal body actions such as atypical head poses and movements, emotions, eye movements, as well as upper and lower limb movements [57, 223].
1.3 Vision-based seizure semiology analysis
@@
There is an increasing interest in leveraging video analysis techniques for the automated analysis of movements associated with motor stereotyped manifestations such as those from Parkinson’s disease, tremors, peripheral neuropathy, Huntington’s chorea, ataxia, dystonia, and epilepsy [39, 162, 166, 111]. Vision-based techniques, in particular, have proven valuable in patient monitoring [195]. While general vision-based motion analyses excel in uncluttered settings, they can be unreliable in noisy environments like epilepsy monitoring and intensive care units. These settings pose unique challenges, including varying lighting conditions throughout recording sessions, environmental occlusions (e.g. bed blankets, head wrapping, hospital gown), and interference from non-subject entities (e.g. clinician, nurse, visitor) [219]. While current state-of-the-art deep learning models exhibit promise in addressing these challenges in epilepsy, their effectiveness is still in the early stages. We can anticipate that more advanced video-based automatic semiologic categorization will become possible as the field progresses. Existing methods struggle to recognize subject-specific semiologic categories of unseen behaviors, yielding modest performance. Achieving fine-grained semiology recognition, crucial for distinguishing the stepwise progression of clinical features, remains a challenge. In addition, ensuring effective explainability to understand model behavior beyond traditional performance indicators is essential, with a particular focus on clinician usability. As these translations of research progress, there is a need for research to facilitate the proper clinical adoption of methods and to build trust between clinical experts and AI frameworks.
1.4 Scope of review and systematic review methodology
@@
As the field is rapidly evolving and researchers are interested in understanding the strengths and pitfalls of deep learning-based solutions for the analysis of seizure semiology, a comprehensive survey is required that provides insights from existing studies and motivates researchers to address limitations and challenges within this field. Conventional motion analysis of seizures, employing classical markerless computer vision methods, was reviewed by Pediaditis et al. [177] and Abbasi et al. [1]. A more detailed examination, considering data-driven solutions for motion capture and action detection, was presented by Ahmedt-Aristizabal et al. [7]. Karacsony et al. [119] provided insights into advancements in video-based clinical in-bed monitoring, particularly within EMUs. These works discuss opportunities to address challenges associated with 3D motion capture, including domain transfer, occlusions, varying viewpoints, and dealing with low resolution footage. A recent study [132] outlined the anticipated workflow of a “complete system” for video analysis of seizure event classification. However, this non-technical discussion overlooked crucial stages.
The application of deep learning approaches to seizure video analysis is still in its nascent stages. The goal of this review is to provide an overview of advances in the field, focusing on work published from 2017 and onwards (the date of our previous review), and highlight ongoing challenges and future directions. The following search string was used in the Google Scholar search engine as an initial search action: (human OR patient) AND (Epilepsy OR epileptic) AND (video OR vision OR computer vision OR action recognition) AND (motion OR movement OR seizure OR semiology OR motor) AND (analysis OR detection OR recognition OR classification) AND (automatic OR machine learning OR deep learning OR artificial intelligence). Case reports and other deep learning applications in Epilepsy including neuroimaging (e.g. MRI, fMRI) and brain activity (e.g. EEG, iEEG) were excluded. Further papers were identified by searching publications by the same authors of initially identified relevant papers, following references, and using the “Cited” function. The results were filtered based on their title and abstract. Case studies reported in this paper are obtained from various peer-reviewed journals, conference proceedings, and open-access repositories published in English. The total number of applications considered in our survey is 35 manuscripts.
1.5 Contribution and organization
In contrast to other reviews that focus on theoretical aspects and applications of action recognition in semiology, our manuscript makes novel contributions, summarized as follows:
@@ 1.5 Contribution and organization
  • We extend our 2016 review of vision-based approaches for semiology analysis [7] by incorporating all relevant studies concerning the adoption of deep learning techniques to enhance the automated detection and quantification of seizures, and/or of individual seizure features captured by video recordings. All deep learning methods adopted by these works are described in detail and categorized according to the motion analysis framework used.
@@
  • Our paper identifies several challenges specific to action recognition models in the context of seizure semiology based on video analysis, which have not been comprehensively addressed in previous reviews. These challenges cover areas such as motion capture of stereotyped movements and domain adaptation, privacy preservation, interpretation of action models, open-set action analysis, and time-evolving or fine-grained action recognition. Existing research efforts in AI and computer vision addressing these challenges are highlighted.
@@
This review is divided into five major sections. Section 2 elucidates concepts and methods in quantifying behavior and behavioral sequences from a neurobiological/neuroscience perspective. Then, the section explores the application of computational neuroethology tools in evaluating these behavioral patterns. In Section 3, we provide a technical overview of human action recognition (HAR) and current approaches for vision-based seizure semiology analysis. Specifically, Section 3.3, introduces skeleton-based HAR approaches, while Section 3.4 focuses on frame-based HAR models. In Section 4, we introduce the current applications of computer vision and deep learning techniques for semiology analysis. Details regarding the existing approaches along with the computer vision techniques are categorized according to the motion framework employed. Section 5 offers guidance on incorporating existing approaches along with the computer vision techniques they employ for clinical use through a multi-stream framework for semiology analysis. Finally, in Section 6, we clarify limitations and offer new perspectives on the current state of semiology-based seizure analysis.
2 Human movement analysis in the epilepsies: from traditional clinical observation to quantitative and computational neuroethology
The evolution of technology has played a pivotal role in shaping our understanding of human behavior, especially in the context of conditions such as motor epilepsies. Computational tools have become integral in uncovering the origins, triggers, and underlying mechanisms of epileptic manifestations, marking a significant advancement in neuroscience through digital and computational technology.
In a recent review [87], we proposed transitioning from traditional clinical observations to quantitative semiology, acknowledging the challenge of unknown neural substrates in earlier eras. This paradigm shift aligns with the systematic evaluation of behavior in animals and humans within the science of ethology [221], involving the creation of dictionaries for detailed behavior descriptions. In the particular case of characterizing semiology, focusing on behavioral features, pioneering studies discussed in [87] have found detailed associations, distinguishing patients with epilepsy from similar clinical entities such as psychogenic non-epileptic seizures (PNES) or cases with neuropsychiatric dysfunctions.
In the progression of this research, the concept of neuroethology emerges as the search for neural substrates of behaviors [77]. Expressions such as the “Grammar of a movement sequence” have been utilized by authors in the field to describe this pursuit. Laboratories initiated the development of ethograms and behavioral glossaries [85]. These protocols involved annotating every behavior and its probabilistic interactions displayed by animals when stimulated, modeling various seizure types [88, 81, 41]. While the term was not explicitly used then, these practices formed the foundation for what we now refer to as quantitative human neuroethology.
Quantitative neuroethology of behavioral sequences served as the natural tool for studies, particularly in epileptic seizures. Initially conducted using animal models [86, 89], these studies were later extended to human patients with both temporal lobe [56, 30] and frontal lobe [31] epileptic seizures. For example, behavioral sequences were reconstructed as flow charts or graphs, with measurements of complexity associated with the expression of the respective semiology [216]. Other researchers have proposed a paradigm shift, viewing semiology as being dynamically produced by interconnected structures, emphasizing the role of specific rhythmic interactions between connected structures within specific networks in clinical expression [159].
An ongoing challenge is the expression of overlapping/redundant substrates of associated interictal neuropsychiatric comorbidities in the behavioral aspects of epilepsy. For example, studies with the concept of autobiographical memories [187] are vital, yet are often excluded in the criteria used to define epileptogenic regions.
Building upon these foundations, we believe that evaluating complex experimental/human behavioral sequences in the context of epilepsy semiology can be significantly enhanced with the deep learning and computer vision approaches discussed in this review. Strengthening this proposal, Krakauer et al. [133] emphasized the importance of behavior studies in neuroscience, advocating for a departure from contemporary reductionistic molecular/cellular biases.
Another challenge and intriguing pathway for advancements in human studies is the adaptation of the concept of quantitative neuroethology to computational neuroscience and modeling. Coined as computational neuroethology [58, 240], this concept focuses on automatically reconstructing complex behavioral fragments, behaviors, or sequences through digital video recording and analysis of individuals or groups of experimental animals. A pioneering program in this field, known as MoSeq2 [238], autonomously collects serial/sequential sets of behaviors. MoSeq2 utilizes a machine learning-based platform to identify behavioral modules (“syllables”) from keypoint data without human supervision.
In conclusion, the fusion of quantitative neuroethology, computational neuroscience, and advanced technologies, including computer vision and deep learning, emerges as a powerful approach to unravel the intricate semiology of epilepsies. This integrated approach not only contributes significantly to diagnosis and treatment strategies, but also signifies a crucial step forward in addressing the challenges posed by this complex neurological disorder.
@@ 3.1 Overview of modern deep learning and human action recognition
Deep learning, a subset of machine learning, enables the creation of computational architectures comprising multiple processing layers. These architectures learn data representations with varying levels of abstraction through non-linear transformations [28]. The primary distinction from traditional machine learning lies in feature engineering. In deep learning algorithms, feature representations are automatically derived from the training data rather than relying on human assumptions as seen in traditional hand-crafted approaches [137]. This proves particularly beneficial in scenarios where understanding the relationship between input and output is limited [15]. Deep learning techniques have become predominant in computer vision, outperforming traditional approaches across various tasks, including human action recognition. Noteworthy deep neural network architectures include convolutional neural networks (CNNs), recurrent neural networks (RNNs), graph neural networks (GNNs) and the Transformer model.
@@
Deep learning methods may also be trained in different ways. In a supervised learning paradigm, training occurs offline as the machine learning processes use the labeled data to learn a mapping from the input to the output, and thus predict output for unforeseen data. However, the scarcity of labeled clinical data and challenges in acquiring additional data make efficient data utilization crucial. To overcome the limitations posed by small datasets, HAR architectures rely on transfer learning. Networks are pre-trained on large-scale datasets [92] which allows the architecture to learn a robust general representation, which can then be applied to extract features and learn a representation specifically for the target domain such as seizure video recordings. However, the generalizability of trained classifiers in subject-independent classification settings is hindered by the considerable variation in behavior across different subjects. A research direction with significant potential is self-supervised learning. In scenarios where no class labels are available, an effective representation could potentially be learned through an end-to-end method entirely in an unsupervised manner, utilizing a large volume of unlabeled data.